# Startup Acquisitions: Acquihires and Talent Hoarding

### v4

@@ 4 Hiring and Layoffs
The adverse shocks come with a commonly known joint probability distribution over “downgrades” for firms, though each firm only observes the realization of its own shock. More specifically, firm $i$ is hit by a shock $S_{i}\in\{D,N\}$, where it is either downgraded from high to low match (if possible) or not affected by the shock. Thus, if a low-match firm is hit by a downgrade, it stays a low-match firm, while a high-match firm turns into a low-match firm. If a firm is not affected by the shock, its match quality stays the same. [^12] Of course, an economy-wide adverse shock may affect more parameters than just the profitability of the high-match firm (for example, it may impact the profitability of startups or firms even in the absence of any acquisitions). We are focusing on the impact on the high-match firms as they will presumably seek to expand, the profitability of which is particularly likely to be impacted by a negative shock to the economy. Let $(S_{1},S_{2})\in\{D,N\}^{2}$ be the profile of shocks hitting the firms, which follows the distribution
@@
where $\gamma\in(0,1)$ is the probability that a firm will be downgraded and $r\in[0,1]$ measures the positive correlation between the firms’ shocks. In particular, for $r=0$ the shocks are independent and for $r=1$ they are perfectly positively correlated.
@@
The presence of talent-hoarding motives always leads to more hiring than in the benchmark. Additionally, provided that $\min\left\{\frac{\lambda_{A}}{\lambda},\frac{1-\lambda}{\lambda}\right\}>(1-r)(1-\gamma)$, talent hoarding also leads to more layoffs than in the benchmark.
@@
After Google announced they would be laying off 12,000 people in 2023 (and following layoffs at other tech companies), a tech commentator John Gruber wrote the following: “There are numerous reasons the tech industry wound up at this layoffpalooza, but I think the main reason is that the biggest companies got caught up in a game where they tried to hire everyone, whether they needed them or not, to keep talent away from competitors and keep talent away from small upstarts (or from founding their own small upstarts). These big companies were just hiring to hire, and now the jig is up.” [^14] From https://daringfireball.net/linked/2023/01/20/google-layoffs, accessed on June 28, 2024. Sam Lessin, the founder of Drop.io mentioned in the introduction, expressed a similar sentiment in an interview with CNBC. [^15] Available at https://www.cnbc.com/video/2024/02/06/tech-companies-were-hoarding-the-talent-says-slow-ventures-sam-lessin.html, accessed on June 28, 2024. Thus, the main economic forces that our model identifies align with industry insiders’ observations.
@@ 5.1 People and Technology
One can verify that for $\tau=0$ the above condition reduces to the condition (1) in the main text. As $\tau$ increases, the threshold $\lambda_{A}(\tau)$ decreases, i.e., the acquisition happens for a larger set of parameters. Hence, a low-match acquirer indeed has an incentive to sell the technology to a high-match competitor because the price compensates her for any decrease in profit due to facing a more efficient competitor. Thus firms do not have an incentive to hoard technology, while the incentive to hoard talent remains. Interestingly, talent hoarding now occurs for a strictly larger set of parameters, as the option to resell technology effectively subsidizes talent hoarding.
@@ 5.2 Partial Acquisitions
6 Discussion and Conclusion
We have presented a simple, yet general, reduced-form model of startup acquisitions. We showed that acquihires may not only reflect firms’ desire to hire talented employees but also be rooted in an incentive to engage in inefficient talent hoarding, thus potentially warranting regulators’ attention. Further, we showed that acquihires may decrease consumer surplus and increase job volatility of acquihired employees, thereby giving further reasons for regulatory scrutiny. Finally, following the inclusion of realistic features into our baseline model such as the startup owning valuable technology that can be resold or the possibility of investing in the startup rather than acquiring it outright, talent hoarding not only persists but may even be exacerbated.
Our baseline model relies on several simplifying assumptions. As we discuss below, the main results hold even if we relax some of those assumptions. The formal analysis and results can be found in Appendix B.
Dominant firm.
Instead of symmetric firms, the market could be characterized by a dominant firm and a challenger. We show that the incentives to hoard talent also emerge in this asymmetric model. Further, under reasonable assumptions about the impact of acquisitions on the profits of the dominant firm and the challenger, the dominant firm is more likely to hoard talent.
Multiple firms.
Here we examine markets with multiple firms that may sequentially try to acquihire the startup. Focusing on the Cournot-Oligopoly setting, we show that in the limit, as the number of firms grows large, no talent hoarding takes place. Intuitively, the profit at risk from a competitor’s acquihire becomes smaller as the number of competitors increases, reducing the incentive to engage in costly talent hoarding. However, the effect of the increase in the number of competitors on the incentive to hoard talent is not necessarily monotonic. The reason for the non-monotonicity is that an increase in the number of competitors, in addition to decreasing profit at risk, also increases the probability that a high-match competitor will acquire the startup. Which effect is stronger for a small number of firms is not clear. Indeed, we show in a parametric example that we may observe more talent hoarding when there are three firms than when there are two.
Uncertain Order of Moves
As noted above, talent hoarding is likely to be a problem whenever firms can identify highly impactful individuals before their competitors do so. However, when a firm becomes aware of a startup with such employees, it may not know whether its competitors are aware of the startup. This extension takes this into account by allowing for an unknown order of moves, where the order of moves is privately drawn by nature with uniform probabilities. We show that talent hoarding still emerges in that setting.
Surplus Sharing
In our baseline model, the firm has all the bargaining power and thus manages to acquire the startup at a price that leaves the entrepreneur with zero surplus. In this extension, we show that the emergence of talent hoarding does not hinge on this very asymmetric distribution of bargaining power. Instead, we show that talent hoarding emerges as long as the entrepreneur does not have too much bargaining power, as otherwise, the acquisition of the startup becomes prohibitively expensive for the firm.
@@ Appendix B Further Extensions
B.1 Dominant Firm
@@
    $\lambda>\lambda_{D}\equiv\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{D}-\underline{\Pi}_{D}^{H}}.$ (B.1)
@@
    $\lambda>\lambda_{C}\equiv\frac{\pi_{E}+\Pi_{C}-\bar{\Pi}_{C}^{L}}{\Pi_{C}-\underline{\Pi}_{C}^{H}}.$ (B.2)
The corollary follows directly from Proposition 1. An interesting question is under which conditions would the dominant firm be more prone to talent hoarding than the challenger firm, i.e., when is $\lambda_{C}>\lambda_{D}$? The following result gives a set of simple sufficient conditions.
@@
If the following two inequalities hold, then $\lambda_{C}>\lambda_{D}$:
@@
$\lambda_{D}\equiv\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{D}-\underline{\Pi}_{D}^{H}},\quad\text{and }\quad\lambda_{C}\equiv\frac{\pi_{E}+\Pi_{C}-\bar{\Pi}_{C}^{L}}{\Pi_{C}-\underline{\Pi}_{C}^{H}}.$
@@
$\lambda_{D}=\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{D}-\underline{\Pi}_{D}^{H}}<\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{C}-\underline{\Pi}_{C}^{H}}<\frac{\pi_{E}+\Pi_{C}-\bar{\Pi}_{C}^{L}}{\Pi_{C}-\underline{\Pi}_{C}^{H}}=\lambda_{C},$
@@ B.2 Multiple Firms
B.3 Uncertain Order of Moves
Proposition B2.
$\lambda\geq\lambda_{A}^{\prime}\equiv\frac{\pi_{E}+\underline{\Pi}^{L}_{F}-\bar{\Pi}_{F}^{L}}{\underline{\Pi}^{L}_{F}-\underline{\Pi}_{F}^{H}}.$ (B.8)
Proof.
Suppose that firm $j$ behaves as suggested in the proposition and that the entrepreneur accepts bids if and only if they are at least $\pi_{E}$. We consider the incentives of firm $i$. Given Assumption 1, it is a dominant strategy for a high-match firm $i$ to make a bid $\pi_{E}$ to do an acquihire whenever it has the chance to do so. Now, consider a low-match firm $i$. Given that firm $j$ would always do an acquihire, firm $i$ knows that it is moving first. Hence, doing nothing yields a payoff of $\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\underline{\Pi}_{F}^{L}$, while doing an acquihire yields $\bar{\Pi}_{F}^{L}-\pi_{E}$. Hence, doing the acquihire is optimal if and only if $\lambda\geq\lambda_{A}^{\prime}$, completing the proof.
B.4 Surplus Sharing
We modify our baseline model to allow for arbitrary degrees of surplus sharing between the entrepreneur and the firm when an acquihire takes place. Thus, firms still move sequentially, but in case of an acquihire the entrepreneur receives a share $\gamma\in[0,1]$ of the surplus. We define surplus here as the difference between the joint payoffs arising from an acquihire and the joint payoffs arising in the case of no acquihire.
Proposition B3.
Talent hoarding can happen if the following condition holds:
$\gamma\leq\frac{\bar{\Pi}_{F}^{L}-\underline{\Pi}_{F}^{H}-\pi_{E}}{\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}}.$
Proof.
We solve the game backward. Consider a high-match firm 2. The surplus resulting from an acquihire is given by $\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}>0$ so that an acquihire takes place and the resulting payoffs for firm 2 and the entrepreneur read $\Pi_{F}+(1-\gamma)(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})$ and $\pi_{E}+\gamma(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})$, respectively. Consider a low-match firm 2. The surplus then reads $\bar{\Pi}_{F}^{L}-\Pi_{F}-\pi_{E}<0$ so that no acquihire takes place.
$\bar{\Pi}_{F}^{H}-(\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\Pi_{F})-((1-\lambda)\pi_{E}+\lambda(\pi_{E}+\gamma(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})))$
$=(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})(1-\lambda\gamma)+\lambda(\Pi_{F}-\underline{\Pi}_{F}^{H})\geq 0,$
$(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})(1-\lambda\gamma)+\lambda(\Pi_{F}-\underline{\Pi}_{F}^{H})+\bar{\Pi}_{F}^{L}-\bar{\Pi}_{F}^{H}.$
$\lambda\geq\lambda_{A,S}\equiv\frac{\pi_{E}+\Pi_{F}-\bar{\Pi}_{F}^{L}}{\Pi_{F}-\underline{\Pi}_{F}^{H}-\gamma(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})},$
for $\Pi_{F}-\underline{\Pi}_{F}^{H}-\gamma(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})>0$. We have $\lambda_{A,S}\leq 1$ whenever the condition stated in the proposition holds.

### v5

@@ 5 Hiring and Layoffs
The adverse shocks come with a commonly known joint probability distribution over “downgrades” for firms, though each firm only observes the realization of its own shock. More specifically, firm $i$ is hit by a shock $S_{i}\in\{D,N\}$, where it is either downgraded from high to low match (if possible) or not affected by the shock. Thus, if a low-match firm is hit by a downgrade, it stays a low-match firm, while a high-match firm turns into a low-match firm. If a firm is not affected by the shock, its match quality remains the same. [^14] Of course, an economy-wide adverse shock may affect more parameters than just the profitability of the high-match firm (for example, it may impact the profitability of startups or firms even in the absence of any acquisitions). We are focusing on the impact on the high-match firms as they will presumably seek to expand, the profitability of which is particularly likely to be impacted by a negative shock to the economy. Let $(S_{1},S_{2})\in\{D,N\}^{2}$ be the profile of shocks hitting the firms, which follows the distribution
@@
where $\gamma\in(0,1)$ is the probability that a firm will be downgraded and $r\in[0,1]$ measures the positive correlation between the firms’ shocks. In particular, for $r=0$ the shocks are independent, and for $r=1$ they are perfectly positively correlated.
@@
The presence of talent-hoarding motives always leads to weakly more hiring than in the benchmark. Additionally, provided that $\min\left\{\frac{\lambda_{A}}{\lambda},\frac{1-\lambda}{\lambda}\right\}>(1-r)(1-\gamma)$, talent hoarding also leads to weakly more layoffs than in the benchmark.
@@
The period towards the end of the COVID pandemic serves as a good example for the effect we aim to capture. Analysts have argued that rapid hiring by tech companies during the pandemic was fueled by a massive positive demand shock tied to work-from-home trends and by historically low interest rates that made borrowing inexpensive. As an adverse shock in the form of employees returning to the office and increased interest rates hit big tech, companies like Google were left with over-expanded workforces, prompting widespread layoff (Griffith, 2022; Irwin-Hunt, 2023). Our model suggests that the size of these layoffs may have been magnified by talent hoarding. Indeed, John Gruber, a tech commentator, wrote the following: “There are numerous reasons the tech industry wound up at this layoffpalooza, but I think the main reason is that the biggest companies got caught up in a game where they tried to hire everyone, whether they needed them or not, to keep talent away from competitors and keep talent away from small upstarts (or from founding their own small upstarts). These big companies were just hiring to hire, and now the jig is up.” (Gruber, 2023) Sam Lessin, the founder of Drop.io mentioned in the introduction, expressed a similar sentiment in an interview (CNBC, 2024). Overall, while the main economic forces that our model identifies are unlikely to capture the complete story behind these layoffs, they align with recent developments and are supported by industry insiders’ observations.
@@ 6.1 People and Technology
One can verify that for $\tau=0$ the above condition reduces to the condition (1) in the baseline model. As $\tau$ increases, the threshold $\lambda_{A}(\tau)$ decreases, i.e., the acquisition happens for a larger set of parameters. Hence, a low-match acquirer indeed has an incentive to sell the technology to a high-match competitor because the price compensates her for any decrease in profit due to facing a more efficient competitor. Thus firms do not have an incentive to hoard technology, while the incentive to hoard talent remains. Interestingly, talent hoarding now occurs for a strictly larger set of parameters, as the option to resell technology effectively subsidizes talent hoarding.
@@ 6.2 Partial Acquisitions
7 Conclusion
We have presented a simple, yet general, reduced-form model of startup acquisitions. We showed that acquihires may not only reflect firms’ desire to hire talented employees but also be rooted in an incentive to engage in inefficient talent hoarding, thus potentially warranting regulators’ attention. Further, we showed that acquihires may decrease consumer surplus and increase job volatility of acquihired employees, thereby giving further reasons for regulatory scrutiny.
Our paper is complimentary to the literature on the acquisition of potential competitors since it shows that welfare-reducing acquisitions can occur even when the acquirer and the target are not active (even potentially) in the same market. Moreover, as we show in Section 6.1, hoarding of talent can be more pernicious than preemptive acquisition of technology or physical assets. This is because, in the absence of property rights, the firms cannot negotiate over the reallocation of talent.
Still, it is not clear from our analysis that regulators always ought to intervene when incumbents acquihire startups. As the discussion surrounding Proposition 2 makes clear, even within our baseline model, restrictive policy can decrease consumer welfare by blocking efficient acquisitions. Moreover, Proposition 2(iii) clarifies that improving consumer welfare can require precise information on the likelihood that the match with an incumbent is high or low, which regulators are not likely to posses. Finally, while our results in Section 4 on the order of moves and surplus sharing show that talent hoarding persists in these settings, they also show that its incidence decreases, further limiting scenarios for which regulatory intervention is desirable. In our view, the balance of benefits and harms of stricter regulatory scrutiny of startup acquisitions remains an open question.
@@ Appendix B Further Extensions
B.1 Simultaneous Moves
First, consider the complete-information situation where the entrepreneur tries to maximize revenue from selling the startup to the two firms. Then, we obtain quite directly from Proposition 1 in Jehiel et al. (1996) the following result for the case of a low-type and a high-type firm.
Proposition B1.
The low-type firm buys the startup at price $p^{L}=\bar{\Pi}_{F}^{L}-\underline{\Pi}_{F}^{H}$ whenever $\bar{\Pi}_{F}^{L}+\underline{\Pi}_{F}^{L}\geq\bar{\Pi}_{F}^{H}+\underline{\Pi}_{F}^{H}$. Otherwise, the high-type firm buys the startup at price $p^{H}=\bar{\Pi}_{F}^{H}-\underline{\Pi}_{F}^{L}$.
The only thing to verify is that these prices exceed $\pi_{E}$. This can be seen from $p^{H}=\bar{\Pi}_{F}^{H}-\underline{\Pi}_{F}^{L}\geq\bar{\Pi}_{F}^{H}-\Pi_{F}>\pi_{E}$ (Assumption 1) and because in case of a (revenue-maximizing) sale to the low-type firm we must have $p^{L}\geq p^{H}$ and thus $p^{L}>\pi_{E}$.
The application of Proposition 2 in Jehiel et al. (1996) for the incomplete-information case is not as straightforward, as the frameworks do not quite match anymore. However, their Application 2 (p. 824) comes quite close and suggests that analagous conditions exist to determine when the low-type or the high-type firm succeeds in buying the startup.
B.2 Uncertain Order of Moves
Proposition B2.
$\lambda\geq\lambda_{A,U}\equiv\frac{\pi_{E}+\underline{\Pi}^{L}_{F}-\bar{\Pi}_{F}^{L}}{\underline{\Pi}^{L}_{F}-\underline{\Pi}_{F}^{H}}.$ (B.1)
Proof.
Suppose that firm $j$ behaves as suggested in the proposition and that the entrepreneur accepts bids if and only if they are at least $\pi_{E}$. We consider the incentives of firm $i$. Given Assumption 1, it is a dominant strategy for a high-match firm $i$ to make a bid $\pi_{E}$ to do an acquihire whenever it has the chance to do so. Now, consider a low-match firm $i$. Given that firm $j$ would always do an acquihire, firm $i$ knows that it is moving first. Hence, doing nothing yields a payoff of $\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\underline{\Pi}_{F}^{L}$, while doing an acquihire yields $\bar{\Pi}_{F}^{L}-\pi_{E}$. Hence, doing the acquihire is optimal if and only if $\lambda\geq\lambda_{A,U}$, completing the proof.
B.3 Emerging Competitor
We consider a scenario in which the two incumbents do not coexist. Specifically, only one incumbent is active, while the other firm may enter with probability $q\in[0,1]$. If the second firm does not enter, and there is no acquihire, the first firm makes a profit $\Pi_{M}$. Acquihiring the startup yields a match-dependent profit of $\bar{\Pi}_{M}^{H}$ or $\bar{\Pi}_{M}^{L}$, where—mirroring Assumption 1—we assume $\bar{\Pi}^{H}_{M}>\Pi_{M}+\pi_{E}>\bar{\Pi}^{L}_{M}$. If the second firm does enter, all profits align with those in the benchmark.
Proposition B3.
Suppose Assumption 1 and $\bar{\Pi}^{H}_{M}>\Pi_{M}+\pi_{E}>\bar{\Pi}^{L}_{M}$ hold, then firm 1’s behavior in any perfect Bayesian equilibrium (PBE) is uniquely specified. Namely, if firm 1 is a high match with the startup, it will pursue an acquihire; if it is a low match, it will pursue an acquihire if and only if
$\lambda\geq\lambda_{A,M}\equiv\frac{\pi_{E}-q(\bar{\Pi}_{F}^{L}-\Pi_{F})-(1-q)(\bar{\Pi}_{M}^{L}-\Pi_{M})}{q(\Pi_{F}-\underline{\Pi}_{F}^{H})}.$ (B.2)
Proof.
Suppose firm 1 has not pursued an acquihire. If the competitor does not emerge, we are done and the monopolist’s profit reads $\Pi_{M}$. If the competitor emerges, it follows from Assumption 1 that, for any belief, firm 2 will pursue an acquihire if and only if it is a high type. Moving to stage 1, firm 1’s belief is given by the prior. A high-match firm 1 will always pursue an acquihire by Assumption 1. A low-match firm 1 will do so whenever $q\bar{\Pi}_{F}^{L}+(1-q)\bar{\Pi}_{M}^{L}-\pi_{E}\geq q(\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\Pi_{F})+(1-q)\Pi_{M}$ or, equivalently, $\lambda\geq\lambda_{A,M}$.
B.4 Surplus Sharing
We modify our baseline model to allow for arbitrary degrees of surplus sharing between the entrepreneur and the firm when an acquihire takes place. Thus, firms still move sequentially, but in case of an acquihire the entrepreneur receives a share $\xi\in[0,1]$ of the surplus. We define surplus here as the difference between the joint payoffs arising from an acquihire and the joint payoffs arising in the case of no acquihire.
Proposition B4.
Talent hoarding can happen if the following condition holds:
$\xi\leq\frac{\bar{\Pi}_{F}^{L}-\underline{\Pi}_{F}^{H}-\pi_{E}}{\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}}.$
Proof.
We solve the game backward. Consider a high-match firm 2. The surplus resulting from an acquihire is given by $\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}>0$ so that an acquihire takes place and the resulting payoffs for firm 2 and the entrepreneur read $\Pi_{F}+(1-\xi)(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})$ and $\pi_{E}+\xi(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})$, respectively. Consider a low-match firm 2. The surplus then reads $\bar{\Pi}_{F}^{L}-\Pi_{F}-\pi_{E}<0$ so that no acquihire takes place.
$\bar{\Pi}_{F}^{H}-(\lambda\underline{\Pi}_{F}^{H}+(1-\lambda)\Pi_{F})-((1-\lambda)\pi_{E}+\lambda(\pi_{E}+\xi(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})))$
$=(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})(1-\lambda\xi)+\lambda(\Pi_{F}-\underline{\Pi}_{F}^{H})\geq 0,$
$(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})(1-\lambda\xi)+\lambda(\Pi_{F}-\underline{\Pi}_{F}^{H})+\bar{\Pi}_{F}^{L}-\bar{\Pi}_{F}^{H}.$
$\lambda\geq\lambda_{A,S}\equiv\frac{\pi_{E}+\Pi_{F}-\bar{\Pi}_{F}^{L}}{\Pi_{F}-\underline{\Pi}_{F}^{H}-\xi(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})},$
for $\Pi_{F}-\underline{\Pi}_{F}^{H}-\xi(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})>0$. We have $\lambda_{A,S}\leq 1$ whenever the condition stated in the proposition holds.
B.5 Dominant Firm
@@
    $\lambda>\lambda_{A,D}\equiv\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{D}-\underline{\Pi}_{D}^{H}}.$ (B.3)
@@
    $\lambda>\lambda_{A,C}\equiv\frac{\pi_{E}+\Pi_{C}-\bar{\Pi}_{C}^{L}}{\Pi_{C}-\underline{\Pi}_{C}^{H}}.$ (B.4)
The corollary follows directly from Proposition 1. An interesting question is under which conditions would the dominant firm be more prone to talent hoarding than the challenger firm, i.e., when is $\lambda_{A,C}>\lambda_{A,D}$? The following result gives a set of simple sufficient conditions.
@@
If the following two inequalities hold, then $\lambda_{A,C}>\lambda_{A,D}$:
@@
$\lambda_{A,D}\equiv\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{D}-\underline{\Pi}_{D}^{H}},\quad\text{and }\quad\lambda_{A,C}\equiv\frac{\pi_{E}+\Pi_{C}-\bar{\Pi}_{C}^{L}}{\Pi_{C}-\underline{\Pi}_{C}^{H}}.$
@@
$\lambda_{A,D}=\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{D}-\underline{\Pi}_{D}^{H}}<\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{C}-\underline{\Pi}_{C}^{H}}<\frac{\pi_{E}+\Pi_{C}-\bar{\Pi}_{C}^{L}}{\Pi_{C}-\underline{\Pi}_{C}^{H}}=\lambda_{A,C},$